# Embracing Imperfection: Simulating Students with Diverse Cognitive Levels Using LLM-based Agents

### v1

@@ Abstract
Large language models (LLMs) are revolutionizing education, with LLM-based agents playing a key role in simulating student behavior. A major challenge in student simulation is modeling the diverse learning patterns of students at various cognitive levels. However, current LLMs, typically trained as “helpful assistants”, target at generating perfect responses. As a result, they struggle to simulate students with diverse cognitive abilities, as they often produce overly advanced answers, missing the natural imperfections that characterize student learning and resulting in unrealistic simulations. To address this issue, we propose a training-free framework for student simulation. We begin by constructing a cognitive prototype for each student using a knowledge graph, which captures their understanding of concepts from past learning records. This prototype is then mapped to new tasks to predict student performance. Next, we simulate student solutions based on these predictions and iteratively refine them using a beam search method to better replicate realistic mistakes. To validate our approach, we construct the Student_100 dataset, consisting of $100$ students working on Python programming and $5,000$ learning records. Experimental results show that our method consistently outperforms baseline models, achieving $100\%$ improvement in simulation accuracy.
@@
Tao Wu1, Jingyuan Chen2†, Wang Lin1, Mengze Li3, Yumeng Zhu2, Ang Li1, Kun Kuang1, Fei Wu1†, 1 College of Computer Science and Technology, Zhejiang University, 2 College of Education, Zhejiang University, 3 Hong Kong University of Science and Technology twu22@zju.edu.cn,    jingyuanchen@zju.edu.cn
@@ 1 Introduction
A natural approach to addressing this challenge is fine-tuning LLMs on error-rich datasets Sonkar et al. (2024). For instance, MalAlgoPy defines 20 common equation transformation errors and trains models on misconception examples to reproduce these mistakes. While such fine-tuning can encourage models to generate errors, it also risks embedding incorrect knowledge, potentially degrading overall performance. Moreover, this method overlooks the personalized nature of student learning, as mistakes should be introduced adaptively based on a student’s cognitive state.
@@ 2.1 LLM-based Education Simulation
Recent research explores using large language models (LLMs) to simulate educational roles, supporting teaching strategy evaluation Markel et al. (2023); Luo et al. (2024) and AI teaching assistant development Yue et al. (2024); Kwon et al. (2024). For instance, Wang et al. (2024); Daheim et al. (2024) use LLMs for personalized error-correction, while Liu et al. (2024b) simulate students with different personalities in ITS. SocraticLMLiu et al. (2024a) models students, teachers, and deans to support Socratic teaching and generate training data. Accurately replicating student behavior, including errors, is vital. Existing methods Sonkar et al. (2024) improve error replication by training on error-rich data, but risk embedding incorrect knowledge. To address this, we propose a training-free pipeline that predicts and refines student behavior simulations for more realistic and natural modeling.
@@ 2.2 Student Cognitive State Analysis
Diagnosing students’ cognitive states and predicting their behavior are key challenges in education research. Existing methods fall into two categories: psychometrics-based and deep learning-based approaches Jiang et al. (2022). Psychometrics-based methods Stout (2007); De La Torre (2009) model the relationship between performance and knowledge state using empirical response functions, while deep learning methods Su et al. (2018); Sun et al. (2022); Dong et al. (2025a, b); Lv et al. (2025); Zhou et al. (2025); Lin et al. (2024) leverage neural networks but rely on implicit, parameterized knowledge, making it difficult to assess mastery of specific concepts Wang et al. (2023). This limitation restricts their effectiveness in student behavior prediction and simulation. To address this, we propose a knowledge graph-based approach to model students’ cognitive prototypes. Our explicit, natural language-driven framework enables more precise behavior prediction and supports solution simulation.
@@ 3 Dataset Curation
The simulation requires datasets with three key features: (1) a stable cognitive state for each student, (2) sequential task-solving records, and (3) detailed annotations, including task statements, solving behavior, and corresponding solutions. However, existing datasets often fail to meet all these criteria. Knowledge tracing datasets Liu et al. (2023) primarily focus on correctness, lacking textual task statements, student behavior, and solutions Hu et al. (2023). In contrast, error diagnosis datasets Wang et al. (2024); Daheim et al. (2024) offer rich textual information but lack annotated task-solving sequences, making it challenging to model students’ cognitive states accurately.
@@ 4.2 Cognitive Prototype Construction
$[v_{i,1},\ v_{i,2},\cdots]=\pi_{node}(d_{i},P_{i})$ (1)
where $[v_{i,1},\ v_{i,2},\cdots]$ denote the extracted concepts.
@@
$[e_{i,1},\ e_{i,2},\cdots]=\pi_{edge}([v_{i,1},\ v_{i,2},\cdots])$ (2)
where $[e_{i,1},\ e_{i,2},\cdots]$ denote the extracted edges.
@@
$[c_{i,1},\ \cdots]=\pi_{local}(d_{i},P_{i},[v_{i,1},\ \cdots])$ (3)
where $[c_{i,1},\ \cdots]$ denote the student’s local cognitive state of each concept.
@@ 4.3 Behavior Prediction
To address this issue, we map the student’s cognitive prototype onto the new task to predict their expected behavior. Specifically, for a given task $t_{j}$, the model $\pi_{desc}$ generates a detailed description of the concepts assessed by the task. This description is compared to the nodes in the knowledge graph to compute similarity scores, and the top-$p$ most relevant knowledge concepts are selected as the reference set, denoted as $[v_{1},v_{2},\cdots,v_{p}]$. Using these $p$ concepts, we identify the past learning record $P_{\hat{j}}\ (1\leq\hat{j}\leq M)$ that includes the largest number of these concepts, which is deemed the most relevant task. By combining the student’s overall cognitive states of these concepts with the context of $P_{\hat{j}}$, the model $\pi_{pred}$ predicts the student’s expected behavior of the given task:
$\hat{b}_{j}=\pi_{pred}(t_{j},[C_{1},C_{2},\cdots,C_{p}],P_{\hat{j}})$ (5)
@@ 6 Conclusion
  • Our method is designed to assess general cognitive proficiency based on a student’s overall performance trends. As a result, it may not always accurately predict occasional careless mistakes, such as typos or incorrect code indentation. However, this challenge is not unique to our approach—baseline methods are also susceptible to such noise. While techniques like causal inference and debiasing could help mitigate the impact of confounding factors, fully addressing this issue falls beyond the scope of our current work. We leave this as a direction for future research and provide a more detailed discussion in the Appendix E.5.
@@ A.2 Dataset Statistics
Figure 7: A sample illustration in our dataset.
@@ C.4 Error Attribution and Traceability
Figure 14: Details of the prompt for Con1 metric.
@@ D.1 Model Details
Figure 15: Details of the prompt for Con2 metric.
D.2 Baseline Details
@@
D.3 Metric Details
Appendix E More Experimental Results and Analysis
E.1 End-to-end Comparison on Student_100
@@ E.3 Results on Java and C++ programming
We compare our method with baselines on these 2 datasets, and the results are shown in Table 8 and Table 9. Experimental results indicate that our method consistently outperforms all baselines on both datasets. This provides strong evidence for the robustness of our framework across different programming subjects, platforms, and student profiles.
@@ E.4 In-Depth Analysis Details
Past Learning Record Volumes Acc Con1 Con2
@@ E.5 Bad Case Analysis
In fact, these occasional careless mistakes pose a challenge not only for our method but also for all baseline approaches, as they are equally susceptible to their influence. We attribute this issue to the statistical nature of existing methods, which assess a student’s cognitive level from a statistical perspective, making them ineffective at handling isolated errors.
@@
Notably, similar challenges have been explored in other domains, such as recommendation systems, where causal debiasing and counterfactual reasoning have been employed to address issues like clickbait Wang et al. (2021a, b, 2022). We believe that effectively handling noise in educational data is an important avenue for future research, and we plan to explore this further in subsequent work.
E.6 Distinctions from Knowledge Tracing
At first glance, our task appears structurally similar to Knowledge Tracing (KT), as both involve modeling student learning processes. However, direct comparison with KT methods is not entirely appropriate due to fundamental differences in objectives and generalization capabilities:

### v2

@@ Abstract
Large language models (LLMs) are revolutionizing education, with LLM-based agents playing a key role in simulating student behavior. A major challenge in student simulation is modeling the diverse learning patterns of students at various cognitive levels. However, current LLMs, typically trained as “helpful assistants”, target at generating perfect responses. As a result, they struggle to simulate students with diverse cognitive abilities, as they often produce overly advanced answers, missing the natural imperfections that characterize student learning and resulting in unrealistic simulations. To address this issue, we propose a training-free framework for student simulation. We begin by constructing a cognitive prototype for each student using a knowledge graph, which captures their understanding of concepts from past learning records. This prototype is then mapped to new tasks to predict student performance. Next, we simulate student solutions based on these predictions and iteratively refine them using a beam search method to better replicate realistic mistakes. To validate our approach, we construct the Student_100 dataset, consisting of $100$ students working on Python programming and $5,000$ learning records. Experimental results show that our method consistently outperforms baseline models, achieving $100\%$ improvement in simulation accuracy and realism. Project page: https://mccartney01.github.io/student_sim.
@@
Tao Wu1, Jingyuan Chen1†, Wang Lin1, Mengze Li2, Yumeng Zhu1, Ang Li1, Kun Kuang1, Fei Wu1† 1 Zhejiang University, 2 Hong Kong University of Science and Technology twu22@zju.edu.cn,    jingyuanchen@zju.edu.cn
@@ 1 Introduction
A natural approach to addressing this challenge is fine-tuning LLMs on error-rich datasets. For instance, MalAlgoPy Sonkar et al. (2024) defines 20 common equation transformation errors and trains models on misconception examples to reproduce these mistakes. While such fine-tuning can encourage models to generate errors, it also risks embedding incorrect knowledge, potentially degrading overall performance. Moreover, this method overlooks the personalized nature of student learning, as mistakes should be introduced adaptively based on a student’s cognitive state.
@@ 2.1 LLM-based Education Simulation
Recent research explores using large language models (LLMs) to simulate educational roles, supporting teaching strategy evaluation Markel et al. (2023); Luo et al. (2024) and AI teaching assistant development Yue et al. (2024); Kwon et al. (2024). For instance, Wang et al. (2024b); Daheim et al. (2024) use LLMs for personalized error-correction, while Liu et al. (2024b) simulate students with different personalities in ITS. SocraticLMLiu et al. (2024a) models students, teachers, and deans to support Socratic teaching and generate training data. Accurately replicating student behavior, including errors, is vital. Existing methods Sonkar et al. (2024) improve error replication by training on error-rich data, but risk embedding incorrect knowledge. To address this, we propose a training-free pipeline that predicts and refines student behavior simulations for more realistic and natural modeling.
@@ 2.2 Student Cognitive State Analysis
Diagnosing students’ cognitive states and predicting their behavior are key challenges in education research. Existing methods fall into two categories: psychometrics-based and deep learning-based approaches Jiang et al. (2022). Psychometrics-based methods Stout (2007); De La Torre (2009) model the relationship between performance and knowledge state using empirical response functions, while deep learning methods Su et al. (2018); Sun et al. (2022); Dong et al. (2025a, b); Lv et al. (2025); Zhou et al. (2025b); Lin et al. (2024) leverage neural networks but rely on implicit, parameterized knowledge, making it difficult to assess mastery of specific concepts Wang et al. (2023b). This limitation restricts their effectiveness in student behavior prediction and simulation. To address this, we propose a knowledge graph-based approach to model students’ cognitive prototypes. Our explicit, natural language-driven framework enables more precise behavior prediction and supports solution simulation.
@@ 3 Dataset Curation
The simulation requires datasets with three key features: (1) a stable cognitive state for each student, (2) sequential task-solving records, and (3) detailed annotations, including task statements, solving behavior, and corresponding solutions. However, existing datasets often fail to meet all these criteria. Knowledge tracing datasets Liu et al. (2023) primarily focus on correctness, lacking textual task statements, student behavior, and solutions Hu et al. (2023). In contrast, error diagnosis datasets Wang et al. (2024b); Daheim et al. (2024) offer rich textual information but lack annotated task-solving sequences, making it challenging to model students’ cognitive states accurately.
@@ 4.2 Cognitive Prototype Construction
$v_{i,1},\ v_{i,2},\cdots]=\pi_{node}(d_{i},P_{i})$ (1)
where $v_{i,1},\ v_{i,2},\cdots]$ denote the extracted concepts.
@@
$e_{i,1},\ e_{i,2},\cdots]=\pi_{edge}([v_{i,1},\ v_{i,2},\cdots])$ (2)
where $e_{i,1},\ e_{i,2},\cdots]$ denote the extracted edges.
@@
$c_{i,1},\ \cdots]=\pi_{local}(d_{i},P_{i},[v_{i,1},\ \cdots])$ (3)
where $c_{i,1},\ \cdots]$ denote the student’s local cognitive state of each concept.
@@ 4.3 Behavior Prediction
To address this issue, we map the student’s cognitive prototype onto the new task to predict their expected behavior. Specifically, for a given task $t_{j}$, the model $\pi_{desc}$ generates a detailed description of the concepts assessed by the task. This description is compared to the nodes in the knowledge graph to compute similarity scores, and the top-$p$ most relevant knowledge concepts are selected as the reference set, denoted as $v_{1},v_{2},\cdots,v_{p}]$. Using these $p$ concepts, we identify the past learning record $P_{\hat{j}}\ (1\leq\hat{j}\leq M)$ that includes the largest number of these concepts, which is deemed the most relevant task. By combining the student’s overall cognitive states of these concepts with the context of $P_{\hat{j}}$, the model $\pi_{pred}$ predicts the student’s expected behavior of the given task:
[…]
@@ 6 Conclusion
  • Our current simulation primarily focuses on textual and behavioral patterns without explicitly incorporating multimodal signals, such as visual or auditory cues Wu et al. (2024); Wang et al. (2024a)that may also influence students’ cognitive processes. While this unimodal setting enables clearer analysis of reasoning behaviors, we acknowledge that a more comprehensive understanding of student performance may benefit from integrating multimodal data. Exploring this dimension constitutes a valuable direction for future work.
@@ A.2 Dataset Statistics
Figure 9: Details of the naive prompt.
@@ C.4 Error Attribution and Traceability
Figure 15: Details of the prompt for Con2 metric.
@@ D.2 Baseline Details
D.3 Metric Details
Appendix E More Experimental Results and Analysis
E.1 End-to-end Comparison on Student_100
@@ E.3 Results on Java and C++ programming
We compare our method with baselines on these 2 datasets, and the results are shown in Table 8 and 9. Experimental results indicate that our method consistently outperforms all baselines on both datasets. This provides strong evidence for the robustness of our framework across different programming subjects, platforms, and student profiles.
@@ E.4 In-Depth Analysis Details
Past Learning
Record Volumes
Acc Con1 Con2
@@ E.5 Bad Case Analysis
In fact, these occasional careless mistakes pose a challenge not only for our method but also for all baseline approaches, as they are equally susceptible to their influence Lin et al. (2023); Wang et al. (2023a). We attribute this issue to the statistical nature of existing methods, which assess a student’s cognitive level from a statistical perspective, making them ineffective at handling isolated errors.
@@
Notably, similar challenges have been explored in other domains, such as recommendation systems Wang et al. (2021a, b, 2022), where causal debiasing and counterfactual reasoning Wang et al. (2025); Yan et al. (2025) have been employed to address issues like clickbait. We believe that effectively handling noise in educational data is an important avenue for future research, and we plan to explore this further in subsequent work.
E.6 Distinctions from Knowledge Tracing
At first glance, our task appears structurally similar to Knowledge Tracing (KT) Zhou et al. (2025a, 2024), as both involve modeling student learning processes. However, direct comparison with KT methods is not entirely appropriate due to fundamental differences in objectives and generalization capabilities: